OpenStudy (anonymous):

Can you help me please : lim cot(2x)/cot(3x) as x approaches to 0

OpenStudy (anonymous):

can u draw the problem please

OpenStudy (anonymous):

@dan815 @Michele_Laino @amistre64

OpenStudy (anonymous):

find the limit of cot(2x)/cot (3x) as x approaches to 0 @israa88

OpenStudy (amistre64):

cos(2x)sin(3x) -------------- sin(2x)cos(3x) we might be able to modify these if need be to get them into a more familiar form

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \tan(3x)\cot(2x)\] I guess you can use L' hopital rule

OpenStudy (amistre64):

sin(a+b) = sin(a)cos(b) + sin(b)cos(a) sin(a-b) = sin(a)cos(b) - sin(b)cos(a) --------------------------------- sin(a+b) + sin(a-b) = 2 sin(a)cos(b) 1/2 (sin(a+b) + sin(a-b)) = sin(a)cos(b)

OpenStudy (amistre64):

sin(3x)cos(2x) ----------- sin(2x)cos(3x) sin(5x) + sin(x) ------------- sin(5x) - sin(x)

OpenStudy (amistre64):

just trying out some manipulations: sin(5x) + sin(x) - sin(x) + sin(x) ---------------------------- sin(5x) - sin(x) 2sin(x) 1 + ------------- sin(5x) - sin(x)

OpenStudy (anonymous):

thank you so much @amistre64

OpenStudy (amistre64):

any of this seem useful to you? its in sin so we might be able to modify it more with like 5x and x stuff

OpenStudy (anonymous):

I'll try it later .. anyway thanks :) @amistre64

OpenStudy (anonymous):

\[\huge \lim_{x \rightarrow 0} \frac{ \cot(2x) }{ \cot(3x) } = \lim_{x \rightarrow 0} \tan(3x)(\cot(2x)) \] You can set it up as 0/0, just another method if you want to check.

OpenStudy (amistre64):

might not have had to go as far as i did, the simple sin cos definintions might be useful cos(2x)sin(3x) -------------- sin(2x)cos(3x) 2x 3x cos(2x)sin(3x) ----------------- 2x 3x sin(2x)cos(3x) 3x cos(2x) ---------- 2x cos(3x)

OpenStudy (anonymous):

That's even better actually :)

OpenStudy (anonymous):

yeah ...

OpenStudy (amistre64):

we got the results now? :)

OpenStudy (amistre64):

cos(2x) and cos(3x) to 1 x/x to 1 leaving ..... tada lol

OpenStudy (anonymous):

:) Voila ! @amistre64